OpenStudy (anonymous):

What number is 5% of 520 show the equation thanks

OpenStudy (calchelp):

5/100 X 520 = (520 x 5) /100 = 2600/100 = 26

OpenStudy (anonymous):

\[X \div520=5\div100\]